The combustion of ethane is given as

C2H6 (g) + 7 O2 (g) --> 4 CO2 (g) + 6 H2O (g)

The rate of formation of carbon dioxide is 0.80 M/s

  1. Calculate the rate of formation of water
  2. Calculate the rate of disappearance of oxygen
  3. Calculate the rate of disappearance of ethane
Answer #1

1) + 1/4 \Delta [CO2] / \Delta t = + 1/6 \Delta [H2O] / \Delta t

\Delta[CO2] / \Delta t = 0.80 M/s  

\Delta[H2O] / \Delta t = 6/4 \Delta [H2O] / \Delta t  

\Delta[H2O] / \Delta t = 6/4 (0.80)

\Delta[H2O] / \Delta t = 1.20 M/s

2)   1/4 \Delta [CO2] / \Delta t = - 1/7 \Delta [O2] / \Delta t

\Delta[O2] / \Delta t = - 7/4  \Delta[CO2] / \Delta t

\Delta[O2] / \Delta t = -7/4 (0.80)

\Delta[O2] / \Delta t = - 1.40 M/s

3)  1/4 \Delta [CO2] / \Delta t = -  \Delta[C2H6] / \Delta t

\Delta[C2H6] / \Delta t = - 1/4 (0.80)

\Delta[C2H6] / \Delta t = - 0.20 M/s